# Real-Time, Population-Based Reconstruction of 3D Bone Models via Very-Low-Dose Protocols

### v1

@@ Abstract
In this study, we collected paired training data consisting of biplanar X-rays (inputs) and 3D bone models (as the ground truth for supervision) manually annotated from corresponding CT volumes. Then, we proposed a new reconstruction framework Semi-Supervised Reconstruction with Knowledge Distillation (SSR-KD) to perform fast and accurate bone model reconstruction from biplanar X-rays. Instead of presenting bone models as mesh or volumes, we regard the bone model as an implicit field [37, 38] (specifically, occupancy field) to improve learning efficiency and achieve better reconstruction quality and higher resolution. Considering the time-consuming labeling, we additionally utilized unlabeled data for training and leveraged semi-supervised learning and knowledge distillation to further improve the reconstruction quality. We quantitatively and qualitatively evaluated our method on a clinical dataset, showing superior reconstruction quality (resolution 256 and average error 1.0 mm) and fast reconstruction speed (30 seconds). More importantly, we further conducted clinical applicability evaluation by performing high tibial osteotomy (HTO) simulation, wherein experienced surgeons and medical professionals evaluated the fitting, stability, and accuracy of surgical guides designed based on 2-view reconstructed and the corresponding CT-based annotated bone models. The results confirmed that surgical guides designed based on bone models reconstructed from biplanar X-rays using our SSR-KD exhibit comparable accuracy to those designed using bone models manually annotated from CT, which validated the feasibility of our proposed algorithm in generating bone models suitable for clinical practice.
@@ Results
We formulated the reconstruction as the estimation of the 3D occupancy field, where 3D points inside the bone surface have an occupancy value of , and points outside the bone surface have an occupancy value of . As shown in Fig. 1b, given anterior-posterior (AP) and right-left (RL) X-ray images as the input, our proposed approach predicts the occupancy values of points in the 3D space to form the occupancy field; then, the surfaces are extracted from the occupancy field using Marching Cubes [1]. The deep neural network was designed to learn a mapping function from a 3D point to a scalar occupancy value. Specifically, as shown in Fig. 7c (also see Reconstruction Networks in the Methods section), the encoder network first extracts the semantic features from input images; then, for a 3D point, the pixel-aligned features are queried from 2D feature maps by its spatial coordinates using bilinear interpolation. Finally, fully connected layers are used to predict the occupancy value. The proposed reconstruction approach achieved an inference time of under 30 seconds, significantly faster than the time required for manually annotating bone models from CT.
@@
Method | Metric | Average | Patella | Femur | Fibula | Tibia |
SSR-KD (ours) | DSC (%) | 90.9 | 87.5 | 96.1 | 84.1 | 96.0 |
HD (mm) | 2.76 | 3.30 | 2.20 | 2.87 | 2.06 | |
ASSD (mm) | 0.94 | 1.12 | 0.80 | 1.05 | 0.77 | |
Baka et al. 2012 [16] (SSM) | ASSD (mm) | - | - | 1.48 | - | - |
Klima et al. 2016 [20] (SSIM) | ASSD (mm) | - | - | 1.18 | - | - |
Kasten et al. 2020 [34] (CNNs) | ASSD (mm) | 1.30 | 1.71 | 1.08 | 1.22 | 1.18 |
@@
For quantitative evaluation, we followed previous works on bone reconstruction [20, 41] to use average symmetric surface distance (ASSD, mm) and Hausdorff distance (HD, mm) to measure the reconstruction error between predicted bone models and ground-truth models (annotated from CT). We also reported the Dice similarity coefficient (DSC, %) based on voxelized 3D bone models. The comprehensive formulation of the aforementioned metrics is provided in the Methods section; refer to the Evaluation of Reconstruction Quality subsection. An independent testing dataset consisting of 40 data pairs was used for evaluation. Quantitative results in Tab. 1 demonstrate that our proposed SSR-KD achieved superior performance than previous studies with an average ASSD below 1 mm. In particular, the surface error (ASSD) of femur and tibia is lower than 0.8 mm, which is very close to the average voxel-wise spacing (0.68 mm) of CT scans. We observed that the shapes of patella and fibula are partially or completely occluded in input views, resulting in higher reconstruction error than femur and tibia.
@@
We conducted a user study to compare the clinical difference between CT-based annotated bone models and 2-view reconstructed bone models. Specifically, we designed a questionnaire from 40 test cases. For each case, we prepared its 2-view X-rays (i.e., AP and RL views), the corresponding CT-based annotated bone models, and the corresponding 2-view reconstructed bone models. Then, each case involved 2 questions – 1.) 2-view X-rays and CT-based annotated bone models rendered from six different viewpoints; 2.) 2-view X-rays and 2-view reconstructed bone models rendered from six different viewpoints. An example of questions is shown in Figure 3a, where neither the case ID nor the version of the bone models was disclosed. A total of 80 questions were collected and randomly shuffled.
For each question, the user was asked to examine the 2-view X-rays and then evaluate the quality of bone models in terms of shape, detail, and clinical significance using four scales (“poor”, “fair”, “good”, and “perfect”). A detailed explanation of four scales is provided in the Methods section; see Evaluation of Reconstruction Quality. We collected evaluation results from 10 experts, including qualified orthopedists, medical school professors, and medical engineers. Each expert evaluated all cases, assessing a total of 80 bone models (two versions per case). Compared to CT-based annotated models (see Fig. 2b), our 2-view reconstructed models obtained comparable scores on all three metrics, which means that experts could hardly identify the difference between CT-based annotated and 2-view reconstructed models given only biplanar X-rays as reference.
@@
We invited surgeons to perform a high tibial osteotomy (HTO) simulation to validate the clinical applicability of patient-specific surgical guides designed based on reconstructed bone models. HTO [42, 43] is a surgical procedure aimed at realigning the knee joint by cutting and reshaping the tibia to reduce pressure on the damaged cartilage, often used to treat knee osteoarthritis or joint deformities. In HTO, surgical guides are designed to fit the patient’s anatomy and ensure precise bone cuts or drill placements [44, 45]. Details of the HTO simulation are illustrated in Fig. 4a(ii)-(iv). For each patient, we designed two versions of surgical guides based on CT-based annotated bone models and our 2-view reconstructed bone models using Materialise 3-matic [46], respectively. Examples of surgical guides are shown in Fig. 4a(i). In addition to 3D printing two versions of surgical guides, we 3D-printed CT-based annotated bone models to replicate the real patient’s anatomy.
During the HTO simulation, the surgeons evaluated each surgical guide using three scores, fitting, stability, and accuracy, to assess how well the surgical guide conforms to the patient’s anatomy, maintains its position during surgery, and ensures precise guidance for bone cuts, which are essential for achieving successful surgical outcomes [47, 48, 49]. The scores were evaluated on a scale of 1 to 6, based on surgeons’ subjective experience. The duration of each HTO operation was also recorded, starting from the placement of the surgical guide on the bone model to the completion of the cut. Detailed explanations on evaluation scores are provided in the Methods section; see Evaluation of Reconstruction Quality. HTO simulation was conducted in a blinded manner for the surgeons, which means that during the procedure, the surgeons did not know which version of surgical guide they were using–whether it was designed based on 2-view reconstructed or CT-based annotated bone models. Surgical guides of five patients were evaluated, and the results are shown in Fig. 4b. On average, 1.) surgeons spent less time with surgical guides designed based on 2-view reconstructed bone models compared to those designed based on CT-based annotated bone models; 2.) surgical guides designed based on 2-view reconstructed bone models achieved comparable scores to those designed based on CT-based annotated bone models. This reveals that the bone models reconstructed by SSR-KD from biplanar X-rays have the potential to be used in surgical guide design during HTO preoperative planning, demonstrating the clinical applicability of our proposed SSR-KD.
@@ Methods
We collected 605 one-leg CT scans covering the entire knee joint with various ages, genders, and scanning ranges of the knee. The dataset specifically comprises 186 female-left, 146 female-right, 142 male-left, and 131 male-right legs, with the field of view varying between 145145108 mm to 375375509 mm, showing the considerable diversity in the data distribution. Among them, 120 CT scans were manually annotated with bone models (patella, femur, fibula, and tibia) by three orthopedic experts, and the remaining 485 CT scans are considered as unlabeled data. CT scans were sampled to the resolution of with the voxel-wise spacing ranging from (0.42, 0.57, 0.57) mm to (1.98, 1.46, 1.46) mm. Annotated bone models were normalized to have coordinates ranging from -1 to +1, and then simplified to have around 8,000 points. We followed previous work [39] to generate anterior-posterior (AP) and right-left (RL) projections from CT scans by digitally reconstructed radiograph (DRR). This process was implemented using TIGRE package with a cone-beam X-ray propagation model. To simulate non-ideal conditions, we also modeled and incorporated statistical quantum and electronic noise. Finally, we have 120 labeled (DRR, CT, bone models) pairs and 485 unlabeled (DRR, CT) pairs. Labeled pairs were split into 70 for training, 10 for validation, and 40 for testing. Therefore, in our experiments, 70 (12.6%) labeled data and 485 (87.4%) unlabeled data were utilized for model training.
@@
Reconstruction from a CT scan (3D). As shown in Fig. 7b, for a 3D CT scan , we employ V-Net [53] as the 3D feature encoder . For a 3D point, the projection is defined as and the implicit function takes the voxel-aligned feature as the input for predicting the 5-dimensional occupancy vector . In practice, is implemented with multi-layer perceptrons, and the channels are 64128256128645.
Reconstruction from 2-view X-ray images (2D). As shown in Fig. 7c, the inputs are 2-view (AP and RL) X-ray projections and . For each input view , we employ an independent (not shared) stacked hourglass network [54] as the 2D feature encoder . The projection function is defined as the same as DRR to project a 3D point to the 2D imaging plane of -view. For a 3D point , we calculate its projected coordinates and then obtain the pixel-aligned feature from -view. In addition, we utilize the distance of to the imaging plane of -view to provide depth information. Finally, pixel-aligned features and depth values are concatenated as . The implicit function takes the concatenated features as the input and predicts the 5-dimensional occupancy vector . In practice, is implemented with multi-layer perceptrons, where the channels are 51410245122561285.
@@
(1.) Training of CT-based reconstruction network . Because CT scans can provide detailed spatial information, 80 labeled data pairs [CT, bone models] are sufficient for training a CT-based network to achieve good performance. During training, 5,000 points are uniformly sampled from the 3D space, and the cross-entropy loss is computed based on the predicted and ground-truth occupancy vectors. The network is trained for 400 epochs with a batch size of 3 and optimized by a momentum gradient descent optimizer with a momentum of 0.98 and an initial learning rate of 0.01 decayed by 0.1 per 100 epochs.
(2.) Training of Xray-based reconstruction network . The training of X-ray-based network is more challenging due to partial occlusion and insufficient labeled data. To address this, we propose a novel framework SSR-KD by leveraging semi-supervised learning and knowledge distillation. The overview of SSR-KD is shown in Fig. 8a. Specifically, for a labeled data pair and a point , the ground-truth occupancy vector is obtained by checking the containment of in annotated bone models. Hence, the predicted is supervised by using cross-entropy (CE) loss:
@@
In practice, the network is trained for 400 epochs with a batch size of 4 and optimized by a momentum gradient optimizer with a momentum of 0.98 and an initial learning rate of 0.01 decay by 0.1 per 100 epochs. , in Eqn. 7 are empirically set to 0.5 and 0.1, respectively. During training, for labeled data, 2,500 points are sampled near the surface of bone models and 2,500 points are uniformly sampled from the 3D space; for unlabeled data, 5,000 points are uniformly sampled from the 3D space.
@@
To further improve the reconstruction quality, we propose to highlight the reconstruction targets by enhancing the intensity of foreground areas (bones) and suppressing the intensity of background areas (soft tissues). Specifically, as shown in Fig. 8b, taking an X-ray image as the input, a segmentation network (U-Net [55]) is used to generate a binary mask (0: background, 1: foreground). Therefore, the enhanced image is calculated by
where is the scaling weight, is element-wise multiplication, and is an all-one matrix of the same size as and . Finally, we concatenate and as the new input for the 2D encoder of X-ray-based reconstruction network. The segmentation network is trained on 300 annotated data pairs [X-ray, mask] with a batch size of 8 for 400 epochs and optimized by a momentum of 0.98 and an initial learning rate of 0.01 decay by 0.1 per 100 epochs. in Eqn. 8 is empirically set to 0.5. Note that the segmentation network is pre-trained and then fixed during the training of X-ray-based reconstruction network.
Evaluation of Reconstruction Quality
As mentioned in the Results section, we evaluate the reconstruction quality of our method from quantitative metrics, visual comparison, and clinical validation. Here, we formally introduce the definition of the above evaluation metrics and scores.
Quantitative Evaluation
@@
Human-Centric Evaluation: User Study
There are four scales (1 - 4: “poor”, “fair”, “good”, and “perfect”) to evaluate the reconstructed bone model in terms of shape, details, and clinical significance. Specifically,
Shape and Details are used to measure the correctness of reconstructed global shape and local details, respectively. A higher shape/details score indicates that the shape/details of the model better match the given images. Four scores represent (1) more than 30%, (2) 10% - 30%, (3) 5% - 10%, and (4) less than 5% of shape/details are not correctly reconstructed.
Clinical Significance is used to assess the significance of reconstructed models in clinical scenarios. A higher score indicates that the reconstructed model has broader clinical applications. Four scores represent the reconstructed model (1) cannot be used for any clinical practice, or can be used for (2) only diagnosis, (3) the planning of osteotomies in certain bones, like high tibial osteotomy (HTO), (4) the planning of most osteotomies, like total knee replacement, etc.
Clinical Applicability Evaluation: HTO Simulation
@@
1.) Fitting: The fitting of a surgical guide refers to how well the device conforms to the anatomical structures of the patient. A well-fitted guide ensures secure placement on the bone surface, which is essential for accurate guidance during surgery. Specifically:
Evaluation: The fitting score measures the percentage of surface contact between the surgical guide and the bone model. Six scales are used to assess this: (1) less than 10% contact, (2) 10% - 30% contact, (3) 30% - 50% contact, (4) 50% - 70% contact, (5) 70% - 90% contact, (6) 90% - 100% contact.
@@
Custom Design: Surgical guides are often designed using preoperative imaging techniques, such as CT or MRI scans, ensuring a precise match to the patient’s anatomy. For example, custom guides have demonstrated significant reductions in maximum deviation from preoperative plans, achieving deviations of only 2 mm compared to 9 mm for manual techniques [47].
2.) Stability: Stability measures how well the surgical guide stays in place during the cutting or drilling process. Stability is critical, as any movement or shifting can result in inaccuracies and compromise the procedure’s success.
Evaluation: Stability is typically assessed by applying force or pressure to the guide and measuring the loss of contact between the guide and the bone. The six scales for stability evaluation are: (1) more than 90% loss of contact, (2) 60% - 90% loss of contact, (3) 30% - 60% loss of contact, (4) 10% to 30% loss of contact, (5) 1% - 10% loss of contact, (6) less than 1% loss of contact.
Structural Integrity: A stable guide must maintain rigidity under surgical forces. For instance, the stiffness of components like base plates has been shown to significantly affect accuracy during drilling operations [48].
Secure Fixation: Features such as alignment pins and holes for securing the guide are critical for maintaining stability throughout the procedure. These are particularly important for moldable surgical guides that require precise placement and stabilization [48].
3.) Accuracy: Accuracy is perhaps the most critical criterion, as it directly impacts surgical outcomes by ensuring that cuts or resections are performed as planned.
Evaluation: The accuracy score measures how well the surgical guide helps the surgeon achieve precise cuts in terms of position and angle. This is evaluated based on the percentage of good vertical and rotational fitting achieved, as well as the quality of guidance provided by the guide: (1) does not provide any, or provided a (2) minimal, (3) satisfactory, (4) acceptable, (5) accurate, (6) precise guide for the surgeon to make precise cuts and has (1) less than 10%, (2) 10% - 30%, (3) 30% - 50%, (4) 50% - 70%, (5) 70% - 90%, (6) more than 90% good vertical and rotational fitting.
Quantitative Measurements: Studies have demonstrated that using custom surgical guides can significantly improve accuracy. For example, in a cadaveric study, custom-guide-assisted techniques resulted in no violations of accepted error thresholds when aiming for cuts within 4 mm of ideal resection lines [47].
Overall, a well-designed surgical guide should have high scores in all three metrics: fitting, stability, and accuracy. These scores are important indicators of the effectiveness of the surgical guide in guiding the surgeon to make accurate cuts during surgery and ensuring the success of the procedure.
@@
We implemented the network design, model training, and model evaluation using PyTorch [56]. Additionally, the primary Python libraries used include Open3D, NumPy, Trimesh, TIGRE, skimage, and SimpleITK. Details of implementation, including network design and training strategies, are provided in the Methods section. The code will be publicly available at https://github.com/xmed-lab/SSR-KD.

### v2

# Real-Time Reconstruction of 3D Bone Models via Very-Low-Dose Protocols

@@ Introduction
In this study, we collected paired training data consisting of biplanar X-rays (inputs) and 3D bone models (as the ground truth for supervision) manually annotated from corresponding CT volumes. Then, we proposed a new reconstruction framework Semi-Supervised Reconstruction with Knowledge Distillation (SSR-KD) to perform fast and accurate bone model reconstruction from biplanar X-rays. A comprehensive comparison of the CT-based solution and our proposed SSR-KD is illustrated in Fig. 1c. Instead of presenting bone models as mesh or volumes, we regard the bone model as an implicit field [36, 37] (specifically, occupancy field) to improve learning efficiency and achieve better reconstruction quality and higher resolution. Considering the time-consuming labeling, we additionally utilized unlabeled data for training and leveraged semi-supervised learning and knowledge distillation to further improve the reconstruction quality. We quantitatively and qualitatively evaluated our method on a clinical dataset, showing superior reconstruction quality (resolution 256 and average error 1.0 mm) and fast reconstruction speed (30 seconds). More importantly, we further conducted clinical applicability evaluation by performing high tibial osteotomy (HTO) simulation, wherein experienced surgeons and medical professionals evaluated the fitting, stability, and accuracy of surgical guides designed based on 2-view reconstructed and the corresponding CT-based annotated bone models. The results confirmed that surgical guides designed based on bone models reconstructed from biplanar X-rays using our SSR-KD exhibit comparable accuracy to those designed using bone models manually annotated from CT, which validated the feasibility of our proposed algorithm in generating bone models suitable for clinical practice.
@@ Results
We formulated the reconstruction as the estimation of the 3D occupancy field, where 3D points inside the bone surface have an occupancy value of , and points outside the bone surface have an occupancy value of . As shown in Fig. 1b, given anterior-posterior (AP) and right-left (RL) X-ray images as the input, our proposed approach predicts the occupancy values of points in the 3D space to form the occupancy field; then, the surfaces are extracted from the occupancy field using Marching Cubes [38]. The deep neural network was designed to learn a mapping function from a 3D point to a scalar occupancy value; see Reconstruction Networks in Methods section. Specifically, the encoder network first extracts the semantic features from input images; then, for a 3D point, the pixel-aligned features are queried from 2D feature maps by its spatial coordinates using bilinear interpolation. Finally, fully connected layers are used to predict the occupancy value. The proposed approach achieved a reconstruction time of approximately 25.2 seconds (network inference: 20.8s, Marching Cubes: 1.5s, post-processing: 2.9s), significantly faster than the time required for manually annotating bone models from CT.
@@
For quantitative evaluation, we followed previous works on bone reconstruction [19, 41] to use average symmetric surface distance (ASSD, mm) and Hausdorff distance (HD, mm) to measure the reconstruction error between predicted bone models and ground-truth models (annotated from CT). We also reported the Dice similarity coefficient (DSC, %) based on voxelized 3D bone models. The comprehensive formulation of the aforementioned metrics is provided in Methods section; refer to the Metrics for Quantitative Evaluation. An independent testing dataset consisting of 40 data pairs was used for evaluation. Quantitative results in Tab. 1 demonstrate that our proposed SSR-KD achieved superior performance than previous studies with an average ASSD below 1 mm. In particular, the surface error (ASSD) of femur and tibia is lower than 0.8 mm, which is very close to the average voxel-wise spacing (0.68 mm) of CT scans. We observed that the shapes of patella and fibula are partially or completely occluded in input views, resulting in higher reconstruction error than femur and tibia.
@@
We conducted a user study to compare the clinical difference between CT-based annotated bone models and 2-view reconstructed bone models. Specifically, we designed a questionnaire from 40 test cases. For each case, we prepared its 2-view X-rays (i.e., AP and RL views), the corresponding CT-based annotated bone models, and the corresponding 2-view reconstructed bone models. Then, each case involved 2 questions – 1.) 2-view X-rays and CT-based annotated bone models rendered from six different viewpoints; 2.) 2-view X-rays and 2-view reconstructed bone models rendered from six different viewpoints. An example of questions is shown in Fig. 3a, where neither the case ID nor the version of the bone models was disclosed. A total of 80 questions were collected and randomly shuffled.
For each question, the user was asked to examine the 2-view X-rays and then evaluate the quality of bone models in terms of shape, detail, and clinical significance using four scales (“poor”, “fair”, “good”, and “perfect”). A detailed explanation of four scales is provided in Methods section; see Metrics for Human-Centric Evaluation. We collected evaluation results from 10 experts, including qualified orthopedists, medical school professors, and medical engineers. Each expert evaluated all cases, assessing a total of 80 bone models (two versions per case). Compared to CT-based annotated models (see Fig. 3b), our 2-view reconstructed models obtained comparable scores on all three metrics, which means that experts could hardly identify the difference between CT-based annotated and 2-view reconstructed models given only biplanar X-rays as reference. Furthermore, the experts’ rating for the clinical significance of the 2-view reconstructed models surpassed a mean score of 3, indicating a consensus on their utility for planning procedures such as HTO.
@@
We invited surgeons to perform a high tibial osteotomy (HTO) simulation to validate the clinical applicability of patient-specific surgical guides designed based on reconstructed bone models. HTO [42, 43] is a surgical procedure aimed at realigning the knee joint by cutting and reshaping the tibia to reduce pressure on the damaged cartilage, often used to treat knee osteoarthritis or joint deformities. In HTO, surgical guides are designed to fit the patient’s anatomy and ensure precise bone cuts or drill placements [44, 45]. Details of the HTO simulation are illustrated in Fig. 4a. For each patient, we designed two versions of surgical guides based on CT-based annotated bone models and our 2-view reconstructed bone models using Materialise 3-matic [46], respectively. Examples of surgical guides are shown in Fig. 4a, labeled as (2) and (3). In addition to 3D printing two versions of surgical guides, we 3D-printed CT-based annotated bone models to replicate the real patient’s anatomy.
During the HTO simulation, the surgeons evaluated each surgical guide using three scores, fitting, stability, and accuracy, to assess how well the surgical guide conforms to the patient’s anatomy, maintains its position during surgery, and ensures precise guidance for bone cuts, which are essential for achieving successful surgical outcomes [47, 48, 49]. The scores were evaluated on a scale of 1 to 6, based on surgeons’ subjective experience. The duration of each HTO operation was also recorded, starting from the placement of the surgical guide on the bone model to the completion of the cut. Detailed explanations on evaluation scores are provided in Methods section; see Metrics for Clinical Application Evaluation. HTO simulation was conducted in a blinded manner for the surgeons, which means that during the procedure, the surgeons did not know which version of surgical guide they were using–whether it was designed based on 2-view reconstructed or CT-based annotated bone models. Surgical guides of five patients were evaluated, and the results are shown in Fig. 4b. On average, 1.) surgeons spent less time with surgical guides designed based on 2-view reconstructed bone models compared to those designed based on CT-based annotated bone models; 2.) surgical guides designed based on 2-view reconstructed bone models achieved comparable scores to those designed based on CT-based annotated bone models. This reveals that the bone models reconstructed by SSR-KD from biplanar X-rays have the potential to be used in surgical guide design during HTO preoperative planning, demonstrating the clinical applicability of our proposed SSR-KD.
@@ Discussion
In our experiments, while the use of Digitally Reconstructed Radiographs (DRRs) provides a controlled environment for validating the proposed reconstruction algorithm, it relies on a simplified imaging physics model. In clinical scenarios, physical nonlinearities such as scatter and beam hardening, along with geometric non-idealities like gantry jitter, are inevitable. These factors introduce complex artifacts (e.g., shading, streaks, and cupping) that can degrade input quality and potentially affect reconstruction accuracy compared to idealized simulations. However, modeling all these complex physical interactions in a simulation environment is computationally intensive and non-trivial. Instead, we plan to address this gap in our future work by collecting a large-scale dataset of real clinical X-rays. This will allow us to further develop and fine-tune the model, ensuring its robustness and applicability in real-world clinical settings.
@@ Methods
We collected 605 one-leg CT scans covering the entire knee joint with various ages, genders, and scanning ranges of the knee. The dataset specifically comprises 186 female-left, 146 female-right, 142 male-left, and 131 male-right legs, with the field of view varying between 145145108 mm to 375375509 mm, showing the considerable diversity in the data distribution. Among them, 120 CT scans were manually annotated with bone models (patella, femur, fibula, and tibia) by three orthopedic experts, and the remaining 485 CT scans are considered as unlabeled data. CT scans were sampled to the resolution of with the voxel-wise spacing ranging from (0.42, 0.57, 0.57) mm to (1.98, 1.46, 1.46) mm. Annotated bone models were normalized to have coordinates ranging from -1 to +1, and then simplified to have around 8,000 points. We followed previous work [39] to generate anterior-posterior (AP) and right-left (RL) projections from CT scans by digitally reconstructed radiograph (DRR). This process was implemented using TIGRE package with a cone-beam X-ray propagation model. To simulate non-ideal conditions, we modeled statistical quantum noise assuming an incident photon flux of photons per pixel, and incorporated electronic noise using a Gaussian distribution (). Finally, we have 120 labeled (DRR, CT, bone models) pairs and 485 unlabeled (DRR, CT) pairs. Labeled pairs were split into 70 for training, 10 for validation, and 40 for testing. Therefore, in our experiments, 70 (12.6%) labeled data and 485 (87.4%) unlabeled data were utilized for model training.
@@
In Fig. 7b, we represent a network for the reconstruction from a CT scan (3D input). For a 3D CT scan , we employ V-Net [53] as the 3D feature encoder . For a 3D point, the projection is defined as and the implicit function takes the voxel-aligned feature as the input for predicting the 5-dimensional occupancy vector . In practice, is implemented with multi-layer perceptrons, and the channels are 64128256128645.
In Fig. 7c, we design a reconstruction for X-rays (2D) as the input. Specifically, the inputs are 2-view (AP and RL) X-ray projections and . For each input view , we employ an independent (not shared) stacked hourglass network [54] as the 2D feature encoder . The projection function is defined as the same as DRR to project a 3D point to the 2D imaging plane of -view. For a 3D point , we calculate its projected coordinates and then obtain the pixel-aligned feature from -view. In addition, we utilize the distance of to the imaging plane of -view to provide depth information. Finally, pixel-aligned features and depth values are concatenated as . The implicit function takes the concatenated features as the input and predicts the 5-dimensional occupancy vector . In practice, is implemented with multi-layer perceptrons, where the channels are 51410245122561285.
@@
In the first step, we train the CT-based reconstruction network () with labeled data, because CT scans can provide detailed spatial information, 80 labeled data pairs [CT, bone models] are sufficient for training a CT-based network to achieve good performance. During training, 5,000 points are uniformly sampled from the 3D space, and the cross-entropy loss is computed based on the predicted and ground-truth occupancy vectors. The network is trained for 400 epochs with a batch size of 3 and optimized by a momentum gradient descent optimizer with a momentum of 0.98 and an initial learning rate of 0.01 decayed by 0.1 per 100 epochs.
In the second step, we train the X-ray-based network (), which is more challenging due to partial occlusion and insufficient labeled data. To address this, we propose a novel framework SSR-KD by leveraging semi-supervised learning and knowledge distillation. The overview of SSR-KD is shown in Fig. 8a. Specifically, for a labeled data pair and a point , the ground-truth occupancy vector is obtained by checking the containment of in annotated bone models. Hence, the predicted is supervised by using cross-entropy (CE) loss:
@@
In practice, the network is trained for 400 epochs with a batch size of 4 and optimized by a momentum gradient optimizer with a momentum of 0.98 and an initial learning rate of 0.01 decay by 0.1 per 100 epochs. , in Eqn. 7 are empirically set to 0.5 and 0.1, respectively. During training, for labeled data, 2,500 points are uniformly sampled from the 3D space, and another 2,500 points are sampled near the bone surface. Specifically, these near-surface points are generated by applying Gaussian noise ( mm) to points on the bone surface. For unlabeled data, 5,000 points are uniformly sampled from the 3D space. During inference, we reconstruct at a resolution of . To mitigate memory constraints, the 16.8 million query points are processed sequentially in mini-batches of 50,000.
@@
To further improve the reconstruction quality, we propose to highlight the reconstruction targets by enhancing the intensity of foreground areas (bones) and suppressing the intensity of background areas (soft tissues). Specifically, as shown in Fig. 8b, taking an X-ray image as the input, a segmentation network (U-Net [55]) is used to generate a binary mask (0: background, 1: foreground). The U-Net design is highly robust to predict high-quality masks, achieving 96.9% segmentation mIoU on our test set. Then, the enhanced image is calculated by
where is the scaling weight, is element-wise multiplication, and is an all-one matrix of the same size as and . Finally, we concatenate and as the new input for the 2D encoder of X-ray-based reconstruction network. This design ensures that raw information is always preserved, preventing data loss from potential segmentation errors, while serves as an additional cue to guide the network’s attention. Empirically, this design proves effective, reducing the overall reconstruction error (ASSD) by 0.12 mm. The segmentation network is trained on 300 annotated data pairs [X-ray, mask] with a batch size of 8 for 400 epochs and optimized by a momentum of 0.98 and an initial learning rate of 0.01 decay by 0.1 per 100 epochs. in Eqn. 8 is empirically set to 0.5. Note that the segmentation network is pre-trained and then fixed during the training of X-ray-based reconstruction network.
Metrics for Quantitative Evaluation
@@
Metrics for Human-Centric Evaluation
There are four scales (1 - 4: “poor”, “fair”, “good”, and “perfect”) to evaluate the reconstructed bone model in terms of shape, details, and clinical significance. Specifically, Shape and Details are used to measure the correctness of reconstructed global shape and local details, respectively. A higher shape/details score indicates that the shape/details of the model better match the given images. Four scores represent (1) more than 30%, (2) 10% - 30%, (3) 5% - 10%, and (4) less than 5% of shape/details are not correctly reconstructed. Clinical Significance is used to assess the significance of reconstructed models in clinical scenarios. A higher score indicates that the reconstructed model has broader clinical applications. Four scores represent the reconstructed model (1) cannot be used for any clinical practice, or can be used for (2) only diagnosis, (3) the planning of osteotomies in certain bones, like high tibial osteotomy (HTO), (4) the planning of most osteotomies, like total knee replacement, etc.
Metrics for Clinical Applicability Evaluation
@@
Surgical guides are often designed using preoperative imaging techniques, such as CT or MRI scans, ensuring a precise match to the patient’s anatomy. For example, custom guides have demonstrated significant reductions in maximum deviation from preoperative plans, achieving deviations of only 2 mm compared to 9 mm for manual techniques [47]. The Fitting score of a surgical guide refers to how well the device conforms to the anatomical structures of the patient. A well-fitted guide ensures secure placement on the bone surface, which is essential for accurate guidance during surgery. Specifically, the fitting score measures the percentage of surface contact between the surgical guide and the bone model. Six scales are used to assess this: (1) less than 10% contact, (2) 10% - 30% contact, (3) 30% - 50% contact, (4) 50% - 70% contact, (5) 70% - 90% contact, (6) 90% - 100% contact.
A stable guide must maintain rigidity under surgical forces. For instance, the stiffness of components like base plates has been shown to significantly affect accuracy during drilling operations [48]. Features such as alignment pins and holes for securing the guide are critical for maintaining stability throughout the procedure. These are particularly important for moldable surgical guides that require precise placement and stabilization [48]. We use Stability score to measure how well the surgical guide stays in place during the cutting or drilling process. Stability is critical, as any movement or shifting can result in inaccuracies and compromise the procedure’s success. Specifically, the stability score is typically assessed by applying force or pressure to the guide and measuring the loss of contact between the guide and the bone. The six scales for stability evaluation are: (1) more than 90% loss of contact, (2) 60% - 90% loss of contact, (3) 30% - 60% loss of contact, (4) 10% to 30% loss of contact, (5) 1% - 10% loss of contact, (6) less than 1% loss of contact.
Studies have demonstrated that using custom surgical guides can significantly improve accuracy. For example, in a cadaveric study, custom-guide-assisted techniques resulted in no violations of accepted error thresholds when aiming for cuts within 4 mm of ideal resection lines [47]. It is important to identify potential sources of error, such as manufacturing tolerances and user errors during assembly or adjustment. For example, manual adjustments in guide length settings can introduce inaccuracies if not carefully calibrated [48, 49]. The Accuracy score is perhaps the most critical criterion, as it directly impacts surgical outcomes by ensuring that cuts or resections are performed as planned. Specifically, the accuracy score measures how well the surgical guide helps the surgeon achieve precise cuts in terms of position and angle. This is evaluated based on the percentage of good vertical and rotational fitting achieved, as well as the quality of guidance provided by the guide: (1) does not provide any, or provided a (2) minimal, (3) satisfactory, (4) acceptable, (5) accurate, (6) precise guide for the surgeon to make precise cuts and has (1) less than 10%, (2) 10% - 30%, (3) 30% - 50%, (4) 50% - 70%, (5) 70% - 90%, (6) more than 90% good vertical and rotational fitting.
Data Availability
The datasets generated and/or analyzed during the current study are not publicly available because this would compromise the patient confidentiality and privacy agreement with the data-providing hospitals, which prohibits any form of public distribution. The minimal dataset that would be necessary to interpret, replicate, and build upon the findings reported in the article is available from the corresponding authors upon reasonable request.
@@
The source code for this study is publicly available on GitHub at https://github.com/xmed-lab/SSR-KD, including the source code for implementing SSR-KD framework, data generation, and experimental analysis. The code is released under the MIT license. We implemented the network design, model training, and model evaluation using PyTorch [56]. Key dependencies include TIGRE for X-ray projection simulation and PyTorch3D for calculating chamfer distance. Other significant libraries used are Open3D, NumPy, Trimesh, skimage, and SimpleITK. The released codebase was tested on a workstation equipped with two NVIDIA GeForce RTX 3090 GPUs (24 GB), Intel Xeon Gold 5218 CPU @ 2.30 GHz, and 128 GB of RAM.